# Revisiting Static Feature-Based Android Malware Detection

### v1

@@ Abstract
The increasing reliance on machine learning (ML) in computer security, particularly for malware classification, has driven significant advancements. However, the replicability and reproducibility of these results are often overlooked, leading to challenges in verifying research findings. This paper highlights critical pitfalls that undermine the validity of ML research in Android malware detection, focusing on dataset and methodological issues. We comprehensively analyze Android malware detection using two datasets and assess offline and continual learning settings with six widely used ML models. Our study reveals that when properly tuned, simpler baseline methods can often outperform more complex models. To address reproducibility challenges, we propose solutions for improving datasets and methodological practices, enabling fairer model comparisons. Additionally, we open-source our code to facilitate malware analysis, making it extensible for new models and datasets. Our paper aims to support future research in Android malware detection and other security domains, enhancing the reliability and reproducibility of published results.
@@ 1 Introduction
Machine learning (ML) has become indispensable in computer security and is extensively used in literature for malware detection [7, 40, 37]. These models learn patterns from data, adapt to new threats, and scale far beyond the capabilities of traditional signature-based systems. However, despite their potential, the design and deployment of ML systems in security applications face significant challenges [6]. Key issues, as observed in prior works [6, 21, 51], include sampling bias, incorrect labels, data snooping, inappropriate threat models, and unsuitable baselines. Additionally, replication studies of these methods are often neglected, leading to claims that cannot be reliably reproduced. This lack of rigor in the design and evaluation process undermines the validity of many reported findings due to poor reproducibility [41]. Addressing these challenges requires a focus on thorough evaluation practices, ensuring that results are consistently reproducible across different environments and datasets.
@@
Figure 1: Comparison of best and worst performing neural networks in Android malware detection using state-of-the-art active learning method from [15]. Models were initialized with 5 different random seeds. The average F1-score over months differs by 10.6%, and the False Negative Rate (FNR) differs by 13.9% between the two models, despite using the same hyperparameters. This highlights reproducibility challenges in machine learning research for Android malware detection.
@@
An example of such reproducibility issues is illustrated in Figure 1. Here, we analyze the reproducibility of the results reported in [15] for Android malware detection using continuous learning. We replicated the experiments using the publicly available source code and dataset [25]. By using the provided best set of hyperparameters for the Drebin-based dataset [7], we conducted the experiments with five different random seeds. We then compared the best and worst-performing models on the test set. The best model achieved an average F1 score of 79.26% over the test months, while the worst model scored 68.65%, indicating a performance difference of 10.6% solely due to variations in random seed initialization. This example highlights how reliably reproducing prior results in Android malware detection can be fraught with challenges, which, if not adequately addressed, can hinder progress in this research field. Furthermore, such issues can lead to disparities between model performance reported in controlled environments and real-world applications, where the stakes are much higher [6].
In this paper, we explore the specific challenges associated with reproducibility and replicability in Android malware analysis, a critical area where these issues are particularly pronounced. By examining key aspects of the dataset and methodology, we shed light on the factors that undermine the credibility of research outcomes in this domain. We identify issues such as data duplicates, inadequate model selection, poor hyperparameter tuning, and flawed evaluation strategies. These contribute to disparities in results obtained under different experimental settings, leading to variability in model performance and thus hurting reproducibility. We systematically discuss how these factors cause irreproducibility and propose ways to mitigate them.
Our study primarily focuses on reproducing the results of a state-of-the-art method for active learning in Android malware detection [15]. However, we expand the scope by considering more diverse experimental settings and addressing general reproducibility issues in Android malware detection. We utilize six different models: Random Forest [13], Support Vector Machine [7], eXtreme Gradient Boosting [14], Multilayer Perceptron [45], Supervised Contrastive Classifier [55], and Hierarchical Contrastive Classifier [15], previously used in Android malware analysis. We analyze both offline learning, where the model is trained once, and active learning, where the model is periodically retrained with a subset of annotated samples. We perform detailed model selection and hyperparameter tuning for all models in both settings. Our analysis shows how data duplicates affect reproducibility, how adequate hyperparameter tuning of baseline models can yield surprising results compared to more complex models, and offers recommendations for enhancing reproducibility in future research. While we focus on Android malware detection, insights from this work can be extended to other domains, such as Windows malware detection [5, 54], and other security applications, like network intrusion detection [21, 41].
Key Contributions
@@
    We identify and discuss key issues and pitfalls in datasets and methodologies in Android malware analysis that lead to poor reproducibility of previously reported results. We propose solutions to address these pitfalls, aiming to improve reproducibility and enable fairer comparisons across different models.
@@
    We implement state-of-the-art machine learning methods for Android malware detection in both offline and active learning settings. Our rigorous study of model performance across diverse experimental settings reveals that simpler baseline methods can often outperform more complex models when properly tuned.
@@
    We open-source our code (to be made public upon publication) for the wider community to facilitate malware analysis. Our code is easily extensible for new models and tasks, supporting future research on Android malware detection and other security tasks to improve the reliability and reproducibility of published results.
@@ 2 Background & Related Work
Table 1: Summary statistics of the datasets used in the study
@@ 2.1 Reproducibility & Replicability
We adhere to ACM’s definitions of reproducibility and replicability, standardized in 2020 [2]. Computational reproducibility refers to the ability of an independent team to achieve a study’s results using the original study’s artifacts [26]. This ensures that the reported findings are valid under specified conditions. Olszewski et al. [41] conducted a longitudinal study on the reproducibility of security papers published in Tier 1 venues (2013-2022), identifying challenges in reproducing results with author-provided artifacts and proposing mitigation strategies.
@@
Our study examines both aspects of machine learning research for Android malware detection. For instance, the analysis shown in Figure 1 underscores reproducibility issues in prior research, even when author-provided artifacts are used, due to unaccounted randomness in published results. Additionally, we address dataset and methodological issues that must be resolved for a study’s claims to be replicable with independently developed artifacts. While previous works have analyzed ML design and datasets in security [6, 21, 51, 47], there is limited research on enhancing the replicability of ML results for malware detection.
In this paper, we use the term reproducibility to generally encompass both computational reproducibility and replicability.
@@ 2.2 Android Malware Detection
Extracted features are input into machine learning models like Random Forests [13], Support Vector Machines [7], and Multilayer Perceptrons [45] to classify apps as malware or benign. The choice of features and models depends on detection objectives and resources. Static analysis is fast and suitable for large-scale detection, whereas dynamic analysis can be more accurate but resource-intensive [3]. This work focuses on static-feature-based machine learning methods due to their prevalence in the literature [30, 42, 58, 8, 15]. However, the methodological flaws identified and recommendations apply to ML using other feature types as well.
Concept Drift:
Malware detection systems encounter challenges due to the evolving nature of malware, leading to concept drift. This drift can arise from new malware families, behavioral changes, evasion attempts, or updates in API semantics. Traditional supervised learning models trained on static datasets may become outdated as new malware types emerge, reducing detection accuracy [55, 17, 42, 53]. These models often fail to recognize emerging variants, resulting in higher false negatives. To address this, online or continuous learning methods can enable models to adapt to new data continuously and stay current with malware trends [30].
Reproducibility of Research in Android Malware Detection Using ML:
Previous studies have assessed issues in reported results for Android malware detection [59, 28], attempted to replicate prior research [18], or addressed limitations in practical environments [22]. However, these studies often rely on single global metrics like AUC or F1-score, lacking nuanced analysis of monthly performance variations, factors influencing reproducibility, and inherent variances in machine learning models due to their stochastic nature. Our work aims to address these gaps by thoroughly analyzing conditions that can lead to poor reproducibility in different experimental settings and how to mitigate them.
@@ 3 Study settings
3.1 Machine Learning
In this study, we consider two distinct machine-learning settings:
@@ 3.1 Machine Learning
    Offline Learning: In this classical supervised setting, the model is trained once on the training dataset and evaluated on the test set. Offline learning for malware detection involves using annotated data from a specific period (e.g., one year) to train a model and evaluate its performance on future unseen samples.
@@
    Active Learning: While offline settings may suffice when the data distribution remains constant, malware detection is susceptible to concept drift. Active learning addresses concept drift by periodically retraining the model [30, 58, 15]. This cost-effective approach focuses on annotating only the samples where the model is least confident, saving annotation resources. Although determining when to retrain remains challenging, a common practice is to retrain the model monthly. This provides a standard evaluation setting for active learning methods independent of drift detection techniques. We adopt this approach in our study.
We treat these settings separately because each requires different model selection methods, such as hyperparameter tuning, as noted in prior work [15]. We aim to analyze the factors influencing replicability in these settings and enhance the reliability of Android malware detection models.
@@ 3.2 Datasets
We utilize the publicly available dataset from [15], which includes two datasets based on static feature types: Drebin [7] and APIGraph [58]. Hereafter, we refer to these as the Drebin [^1] This dataset is collected from and referred to as ”Androzoo”[1] in  [15] but since it uses the same feature sets as original Drebin dataset [7], we refer to it as Drebin. and APIGraph datasets. Drebin uses eight different sets of features, including access to hardware components, requested permissions, app components, intents, usage of restricted API calls, used permissions, suspicious API calls, and network addresses. APIGraph employs API semantics to cluster similar APIs, significantly reducing the feature space. The resulting feature set is binary for both datasets, indicating whether a feature is present or absent for a specific sample.
The Drebin dataset contains applications collected from 2019 to 2021, while the APIGraph dataset includes applications from 2012 to 2018. Both datasets address temporal and spatial biases present in malware datasets [6, 42]. They consist of approximately 90% benign and 10% malware applications collected from each month, reflecting their real-world ratios. Samples are ordered and approximately evenly distributed over their respective periods. Table 1 provides a summary of the datasets.
Following [15], we use data from the first year as the training set, the next six months for validation, and the remaining months for testing. For offline, only the training set data is used to train the model. In the active learning setup, this data is used to train the initial model, which is then periodically retrained using data from subsequent months from the validation and test set. The validation set is used to tune the model hyperparameters in both experimental settings.
@@
(b) Drebin: Active learning deduplication
@@
(d) APIGraph: Active learning deduplication
@@
4 Pitfalls
In this section, we examine the challenges in the existing literature on malware detection using machine learning. We focus on dataset and methodological issues that impact the reproducibility of reported results.
4.1 Data Duplicates
Issues of duplicates in Android malware datasets have been reported in prior studies [28, 59]. The study by Irolla et al. [28] analyzed duplicates in the original Drebin dataset [7] using opcode sequences and found that 50.65% of malware samples were unique, with the rest being duplicates. They demonstrated that removing duplicates can alter the performance rankings of different machine learning models. Zhao et al. [59] examined duplicates in four malware datasets using APK, DEX, opcode sequences, and API calls to identify them. They also evaluated the impact of duplicates on multiple machine-learning models.
Our study differs from these works in two key ways. First, neither study provided a time-aware analysis of duplicates’ effects on malware detection models. Time-aware metrics are recommended for malware detection [42], and results based only on global metrics (such as F1-score) may not accurately reflect the situation, as the level of duplicates can vary monthly. We will demonstrate this variability later. Second, we evaluate the impact of duplicates in both offline and active learning settings, whereas previous studies focused only on offline learning. These settings require different methods to address duplicates, which we discuss later.
@@ 4.1.1 Defining Duplicates
We define two samples as duplicates if their feature sets are identical after the feature extraction process, meaning they are presented as identical inputs to the machine learning model. Specifically, two samples $X$ and $Z$ are considered duplicates if $x_{1}=z_{1},x_{2}=z_{2},\ldots,x_{n}=z_{n}$ for an $n$-dimensional feature vector:
@@
This definition ensures that duplicate analysis is independent of how features are derived, unlike previous studies [28, 59]. It is important to note that this definition does not include "near-duplicates," such as samples from the same malware family that may have similar behavior but slightly different feature sets due to changes in benign code.
The presence of duplicate samples can bias machine learning models by causing specific training samples to be redundant and skewing performance metrics on the test set [4]. If a sample appears in both the training and test sets, it can artificially inflate test set performance, thus failing to reflect the model’s true generalization abilities accurately. Therefore, it is crucial to consider duplicates when evaluating machine learning models.
@@ 4.1.2 Deduplication
The process of deduplication ensures that the dataset contains unique samples. We define two types of deduplication based on different experimental settings in our study:
Offline Learning Deduplication: In offline learning, the model is trained once on a training set and evaluated on validation and test sets. Deduplication involves selecting unique samples within each data split. The model is evaluated monthly, requiring two levels of deduplication. First, we remove samples that appear in earlier splits. For the validation set, we remove duplicates from the training set. We remove duplicates from the training and validation sets for the test set. Within each split, we keep only the earliest occurrence of a sample. This ensures that all samples across the training, validation, and test sets are unique, retaining the earliest samples based on appearance time.
Active Learning Deduplication: When retraining the model monthly, the definition of duplicates changes due to continuous model updates, which can affect behavior, such as due to catastrophic forgetting [44]. Following [15], we assume that training data is always available. We remove duplicate samples from the training split, which is similar to offline learning. For the validation and test sets, duplicates are removed only within the same month, not across different months. This approach allows duplicates to appear in multiple months of the test set but ensures each month contains unique instances. This approach is essential for active learning as the model continuously evolves. If a sample appears in two months, the model should still recognize it accurately. By maintaining unique samples each month, we assess the model’s performance fairly and prevent duplicates from affecting retraining.

### v2

# R+R: Revisiting Static Feature-Based Android Malware Detection using Machine Learning

@@ Abstract
Static feature-based Android malware detection using machine learning (ML) remains critical due to its scalability and efficiency. However, existing approaches often overlook security-critical reproducibility concerns, such as dataset duplication, inadequate hyperparameter tuning, and variance from random initialization. This can significantly compromise the practical effectiveness of these systems. In this paper, we systematically investigate these challenges by proposing a more rigorous methodology for model selection and evaluation. Using two widely used datasets, Drebin and APIGraph, we evaluate six ML models of varying complexity under both offline and continuous active learning settings. Our analysis demonstrates that, contrary to popular belief, well-tuned, simpler models, particularly tree-based methods like XGBoost, consistently outperform more complex neural networks, especially when duplicates are removed. To promote transparency and reproducibility, we open-source our codebase, which is extensible for integrating new models and datasets, facilitating reproducible security research.
@@ 1 Introduction
The proliferation of Android malware continues to pose significant security threats, making timely detection and mitigation critical. Machine learning (ML), particularly leveraging static analysis features, remains widely adopted in security practice due to its scalability, speed, and suitability for rapidly screening large numbers of applications [1, 2, 3, 4]. ML models learn patterns from data, adapt to new threats, and scale far beyond the capabilities of traditional signature-based systems. However, recent research underscores critical limitations in their reliability, reproducibility, and security assurance [5], leading to claims that cannot be reliably corroborated. Key issues, as observed in prior works [5, 6, 7], include sampling bias, incorrect labels, data snooping, inappropriate threat models, and unsuitable baselines. This lack of rigor in the design and evaluation process undermines the validity of many reported findings  [8]. Such discrepancies seriously question the reliability of ML-based malware detection when deployed in operational security settings.
@@
Figure 1: Performance comparison of the best and worst performing neural network models for Android malware detection using the state-of-the-art continuous active learning method from [9]. Top: F1-score. Bottom: False Negative Rate (FNR). Models were initialized with 5 different random seeds. The average F1-score over months differs by 10.6%, and the False Negative Rate (FNR) differs by 13.9% between the two models, despite using the same hyperparameters. This highlights reproducibility challenges in machine learning research for Android malware detection.
@@
Motivation. We showcase an example that highlights how reliably reproducing prior results in Android malware detection can be fraught with challenges, which, if not adequately addressed, can hinder progress in this research field. See Figure 1, where we examine the reproducibility of results reported by [9] for Android malware detection using continuous active learning. Reproduced experiments include artifacts provided by the authors, including code and datasets, and the author-provided best set of hyperparameters for the Drebin-based dataset [1]. We used five different random seeds and then compared the best and worst-performing models on the test set. Best model achieved an average F1 score of 79.26% over the test months, while the worst model scored 68.65%, indicating a performance difference of 10.6% solely due to variations in random seed initialization.
In this paper, we study these security-critical evaluation issues by systematically studying Android malware detection methodologies. We examine three prevalent reproducibility pitfalls, including (1) data duplication, which artificially inflates model performance; (2) inadequate hyperparameter tuning, causing unfair baseline comparisons; and (3) performance instability due to variance from random model initialization. We systematically discuss how these factors cause irreproducibility and propose ways to mitigate them. We use two widely-recognized Android malware datasets, Drebin [1] and APIGraph [10], and six different models, including Random Forest [11], Support Vector Machine [1], eXtreme Gradient Boosting [12], Multilayer Perceptron [13], Supervised Contrastive Classifier [14], and Hierarchical Contrastive Classifier [9], previously used in Android malware analysis. We conduct comprehensive experiments under realistic offline learning, where the model is trained once, and continuous active learning, where the model is periodically retrained with a subset of annotated samples. We perform detailed model selection and hyperparameter tuning for all models in both settings. While our work builds on recent advances in continuous active learning [9], we broaden the evaluation scope to include diverse settings and emphasize reproducibility in Android malware detection. Contrary to the assumption that complex neural architectures outperform simpler ones, our experiments demonstrate that well-tuned tree-based models, particularly XGBoost, often exhibit greater robustness after duplicate samples are removed. These results underscore the importance of rigorous and reproducible evaluation over architectural sophistication.
Key Contributions:
@@
    We systematically identify critical reproducibility issues in datasets and evaluation methods commonly used in Android malware detection, and propose rigorous methodological solutions to improve result reliability and enable fairer comparisons across models.
@@
    Through comprehensive experiments on two standard datasets in both offline and continuous active learning scenarios, we demonstrate that simpler, well-tuned models (e.g., XGBoost) consistently outperform complex neural networks, challenging common assumptions in security literature.
@@
    We release our extensively validated, extensible evaluation framework as open-source [^1] https://github.com/xashru/maldetect, facilitating transparent, reproducible, and security-focused malware detection research.
@@ 2 Background & Related Work
TABLE I: Summary statistics of the datasets used in the study
@@ 2.1 Reproducibility & Replicability
We adhere to ACM’s definitions of reproducibility and replicability, standardized in 2020 [15]. Computational reproducibility refers to the ability of an independent team to achieve a study’s results using the original study’s artifacts [16]. This ensures that the reported findings are valid under specified conditions. [8] conducted a longitudinal study on the reproducibility of security papers published in Tier 1 venues (2013-2022), identifying challenges in reproducing results with author-provided artifacts and proposing mitigation strategies.
@@
Our study examines both aspects of machine learning research for Android malware detection. For instance, the analysis shown in Figure 1 underscores reproducibility issues in prior research, even when author-provided artifacts are used, due to unaccounted variability in published results. Additionally, we discuss pitfalls that must be addressed for a study’s claims to be replicable with independently developed artifacts. In this paper, we use the term reproducibility to broadly refer to both computational reproducibility and replicability.
@@ 2.2 Android Malware Detection
Extracted features are input into machine learning models like Random Forests [11], Support Vector Machines [1], and Multilayer Perceptrons [13] to classify apps as malware or benign. The choice of features and models depends on detection objectives and resources. Static analysis is fast and suitable for large-scale detection, whereas dynamic analysis can be more accurate but resource-intensive [23]. This work focuses on static-feature-based machine learning methods due to their prevalence in the literature [24, 25, 10, 26, 9].
Concept Drift: Malware detection systems encounter challenges due to the evolving nature of malware, leading to concept drift. This drift can arise from new malware families, behavioral changes, evasion attempts, or updates in API semantics. Traditional supervised learning models trained on static datasets may become outdated as new malware types emerge, reducing detection accuracy [14, 27, 25, 28]. These models often fail to recognize emerging variants, resulting in higher false negatives. To address this, continuous learning methods can enable models to adapt to new data continuously and stay current with malware trends [24, 9].
Reproducibility of Research in Android Malware Detection Using ML: Previous studies have assessed issues in reported results for Android malware detection [29, 30], attempted to replicate prior research [31], or addressed limitations in practical environments [32]. However, these studies do not consider the continuous learning setup, often rely on single global metrics like AUC or F1-score, lacking nuanced analysis of monthly performance variations, factors influencing reproducibility, and inherent variances in machine learning models due to their stochastic nature. Our work aims to address these gaps by analyzing conditions that can lead to poor reproducibility in different experimental settings and how to mitigate them.
@@ 3 Study settings
3.1 Machine Learning Scenarios
We evaluate Android malware detection models under 2 distinct ML scenarios relevant to practical security deployment:
@@ 3.1 Machine Learning Scenarios
    Offline Learning: Models are trained once on the historical dataset and then evaluated on unseen future data, reflecting static deployment scenarios. This involves using annotated data from a specific period (e.g., one year) to train a model.
@@
    Continuous Active Learning: To address the evolving nature of malware (concept drift), we periodically retrain the model [24, 10, 9] using only the samples where the model is least confident, saving annotation resources. We adopt monthly retraining intervals, consistent with established literature.
We treat these settings separately because each requires different model selection methods, such as hyperparameter tuning, as noted in prior work [9].
@@ 3.2 Datasets
We utilize the publicly available dataset from [9], which includes two datasets based on static feature types: Drebin [1] and APIGraph [10]. For brevity, we refer to them as the Drebin and APIGraph datasets, although the Drebin variant used here was collected between 2019 and 2021, rather than the original 2010–2012 corpus [1]. This ensures comparability with the evaluation setup of [9]. Both datasets have been extensively applied across various contexts, including explainable ML [33], concept drift analysis [32, 26], obfuscation studies [34], and semi-supervised learning [35, 36], reflecting their central role in the literature. However, despite their broad use, prior studies have not adequately examined the inherent reproducibility challenges in these datasets, which complicates efforts to replicate results in new environments. Addressing these issues provides the primary motivation for our work.
Drebin uses eight different sets of features, including access to hardware components, requested permissions, app components, intents, usage of restricted API calls, used permissions, suspicious API calls, and network addresses. APIGraph employs API semantics to cluster similar APIs, significantly reducing the feature space. The resulting feature set is binary for both datasets, indicating whether a feature is present or absent for a specific sample.
The Drebin dataset contains applications collected from 2019 to 2021, while the APIGraph dataset includes applications from 2012 to 2018. Both datasets address temporal and spatial biases present in malware datasets [5, 25]. They consist of approximately 90% benign and 10% malware applications collected from each month, reflecting their real-world ratios. Samples are ordered and approximately evenly distributed over their respective periods. Table I provides a summary of the datasets.
Following [9], we use data from the first year as the training set, the next six months for validation, and the remaining months for testing. For offline learning, only the training set data is used to train the model. In the continuous active learning setup, this data is used to train the initial model, which is then periodically retrained using data from subsequent months from the validation and test set. The validation set is used to tune the model hyperparameters in both experimental settings.
@@
(b) Drebin: Continuous active learning deduplication
@@
(d) APIGraph: Continuous active learning deduplication
@@
4 Pitfalls in Reproducibility and Evaluation
We critically examine key methodological and dataset-related pitfalls prevalent in existing Android malware detection research. These pitfalls can severely undermine reproducibility and practical security effectiveness, and we propose systematic strategies to mitigate their impact.
4.1 Dataset Duplication
Issues of duplicates in Android malware datasets have been reported in prior studies [30, 29]. The study by [30] analyzed duplicates in the original Drebin dataset [1] using opcode sequences and found that 50.65% of malware samples were unique, with the rest being duplicates. They demonstrated that removing duplicates can alter the performance rankings of different machine learning models. [29] examined duplicates in four malware datasets using APK, DEX, opcode sequences, and API calls to identify them. They also evaluated the impact of duplicates on multiple machine-learning models.
Our study differs from prior works in key ways. First, we are the first to address the reproducibility of reported results in this context. Second, we provide a time-aware analysis of duplicates’ effects on malware detection models, as global metrics (e.g., F1-score) may miss monthly variations [25]. Lastly, unlike prior studies focused only on offline learning, we assess duplicates’ impact in offline and continuous learning settings, which require different deduplication approaches, as discussed later.
@@ 4.1.1 Defining Duplicates
We define duplicates based strictly on feature-vector equality, meaning two samples are duplicates if their extracted feature vectors are identical. Specifically, two samples $X$ and $Z$ are considered duplicates if $x_{1}=z_{1},x_{2}=z_{2},\ldots,x_{n}=z_{n}$ for an $n$-dimensional feature vector:
@@
This definition directly reflects the model’s perspective, differing from prior approaches relying on raw or opcode-based duplication checks [30, 29]. Notably, despite using this feature-space definition, we observe a similar percentage of duplicates in the datasets as reported in prior studies. Also, detecting accurate duplicates is particularly challenging in malware analysis, which we further discuss in Section 10.
@@ 4.1.2 Deduplication
The process of deduplication ensures that the dataset contains unique samples. We establish rigorous deduplication protocols based on different experimental settings in our study:
Offline Learning Deduplication: Here, the model is trained once on a training set and evaluated on validation and test sets. Deduplication involves selecting unique samples within and across data splits. Since the model is evaluated monthly, two levels of deduplication are required. First, we remove samples that appear in earlier data splits: duplicates in the validation set are removed if present in the training set, while the test set excludes duplicates from both the training and validation sets. Second, within each split, only the earliest occurrence of a sample is retained based on its appearance time.
Continuous Active Learning Deduplication: Here, the definition of duplicates evolves due to monthly model retraining and changes in behavior, such as catastrophic forgetting [37]. Duplicate samples are first removed from the training split, following the same approach as offline learning. For the validation and test sets, duplicates are removed only within the same month, not across different months. This allows duplicates to appear in multiple months of the test set while ensuring that each month contains unique instances. This approach is crucial for continuous learning, as periodic retraining requires the model to recognize recurring samples accurately.